the length of a rectangle is one unit more than it's width. if the area is 30 square units, find the demnsions of the rectangle. translate the problem into an equation.
Answer by supriyo_bol(1)

Let the width be = L unit
Length = (L+1) unit
Area= L(L+1)
Therefore
L(L+1)=30
L2+L-30=0